Circular No. 6587

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) R. West, European Southern Observatory; and M. Kidger, Instituto de

Astrofísica de Canarias, report on behalf of the European Comet Hale-Bopp Team: “High-resolution observations taken with the 1-m Jacobus Kapteyn Telescope of the Observatorio del Roque de los Muchachos by J. Licandro (Instituto de Astrofísica de Canarias) and P. Rudd (Royal Greenwich Observatory) on Mar. 6.2 UT broadly confirm the observations of Birkle and Böhnhardt reported on IAUC 6583. Observations were taken through a CN and a blue continuum filter, with Laplacian filtering applied to the reduced images. The images reveal both spiral-jet and arc structures. The spiral jet is seen in both filters and leaves the nucleus at p.a. ∼ 200^{o}, turning through a right angle at ∼ 3′′ from the nucleus and continuing in p.a. 110^{o}. A bright shell is also seen in both filters 20′′ to the southwest of the nucleus, which corresponds to a recently separated jet. Two diffuse arc structures are also seen to the north and east of the nucleus. The northern arc is only seen in the CN images, centered at 18′′ from the nucleus, and it extends from p.a. 90^{o} to 270^{o} through the north, being clearly elliptical, with the major axis in p.a. 340^{o}. A second and fainter arc can be seen only in the continuum-subtracted CN image and extends from p.a. 30^{o} to 120^{o}, centered ∼ 27′′ from the nucleus. The CN and continuum images can be

found at http://www.iac.es/comet/ima97/mar5ifin.jpg. Assuming, as

on IAUC 6583, that the two arcs are ejected by the same active spot and that the nucleus rotates with an 11.47-hr period (Lecacheaux et al., IAUC 6560; Jorda et al., IAUC 6583), we obtain an expansion velocity for CN of 1.3 km/s for a 60^{o} rotation between arc emissions, very close to the velocity calculated by Birkle and Böhnhardt.”

J. E. Wink, Institut de Radio Astronomie Millimètrique (IRAM), Grenoble, reports on behalf of the continuum and molecular-lines observing teams (headed by W. J. Altenhoff, Max Planck Institut für Radioastronomie; and D. Bockelée-Morvan, Observatoire de Paris, respectively): “We have detected the millimeter-wave continuum of C/1995 O1 with the IRAM Plateau de Bure Interferometer at 3 mm (resolution 3′′.5) and at 1 mm (resolution 1′′.5) on Mar. 9 and 11. On Mar. 11.4 UT, the flux density of the unresolved nucleus was 24 ± 1 mJy at 229 GHz. This flux can be accounted for by thermal radiation at blackbody temperature 380 K from a 1600-km^{2} area, which corresponds to a sphere of diameter 45 km. These observations were made simultaneously with molecular-line observations, which detected cometary emissions centered at the continuum position.”

1997 March 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green